OpenStudy (anonymous):

Paula receives a salary of $350 per week plus 2% on sales up to $5000 and 3% on the remaining amount over $5000. Paula averages $8550 in sales per week. What would be her expected total earnings for the week?

OpenStudy (aripotta):

so she gets 2% for the first $5000 and 3% on what's remaining 8550 - 5000 = 3550 350 + (5000 x 0.02) + (3550 x 0.03)

OpenStudy (anonymous):

350 + (5000 x 0.02) + (3550 x 0.03)

OpenStudy (anonymous):

well did you ues a culalater and then it will work

OpenStudy (aripotta):

2415 isn't even the answer

OpenStudy (anonymous):

okay so this is my answer 350+100 +106.5 = 556.5

OpenStudy (aripotta):

yes, that's right magan :)

OpenStudy (anonymous):

1.02 (5000) + 1.03 (3550) = 8756.50

OpenStudy (anonymous):

Okay thx alot.

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

@magan I removed all of my responses on this one. There is a question as to wether or not to include the base salary of $350 per week.